I don't know how to turn this into an equation.
If someone works 36 hours per week they make 47,073 per year. What is their annual salary if they work 32 hours per week?

Answer by checkley71(8403)

32/36=X/47,073 OR 36X=32*47,073 OR 36X=1,506,336OR X=1,506,336/36 OR
X=41,842.67 ANNUAL SALARY FOR 32HOUR WEEK.